2 2.1 2.2 2.3 2.4 2.5 2.6 2.7 2.8 1980  1990  2000  2010 Bell parameter S  [?, ?, ?] [?] [?] [?]  [?] [?] [?]  [?] [?] [?]  [?]  2.8 2.81 2.82 2.83 2.84 2.85 2013  2014  2015 Bell parameter S  [?] [?] [?] Maxumum violation [?]  Maxumum violation Year  Year [?]

Figure 1: Since the first experimental violation of a Bell’s inequality in 1982 [?], there have been several efforts to violate the inequality with a variety of systems. Photonic systems (Green circles) have, so far been the most successful. Systems comprising of atoms, ions, and atoms and photons (Blue diamonds) have also been studied. Bell violations have also been demonstrated with Josephson junctions (Pink square) and NV centers in diamond (Orange triangle).